# MM-Sonate: Multimodal Controllable Audio-Video Generation with Zero-Shot Voice Cloning]Kling Team, Kuaishou Technology

### v1

@@ 6.3 Evaluation Metrics
  • Video Quality: We report Motion Score (MS) based on RAFT [53] optical flow to quantify dynamic realism. Aesthetic Score (AS) is a composite metric derived from MANIQA [54], aesthetic-predictor-v2-5 [55], and Musiq [56] to assess visual fidelity. Identity preservation is measured by ID Consistency (ID), computed as the mean DINOV3 [57] feature similarity between reference images and generated frames.
  • Audio Quality: We measure the distributional distance using Fréchet Distance (FD) and KL Divergence on mel-spectrogram features extracted via PANNs [58]. Semantic alignment between audio and text is evaluated using the CLAP score [59].
  • Speech & Voice Cloning: Speech intelligibility is quantified by the Word Error Rate (WER), derived from Whisper-large-v3 [60] transcriptions. To assess speaker identity preservation, we employ specific metrics depending on the evaluation context. For the main comparative benchmark against state-of-the-art TTS models (Table 2), we follow the SeedTTS[61] evaluation protocol and report SIM-o, calculated using the WavLM-large [39] speaker verification model. In our ablation studies regarding negative conditioning strategies (Fig. 3 and 4), we utilize the Resemblyzer encoder***https://github.com/resemble-ai/Resemblyzer to measure Speaker Consistency.
  • Audio-Visual Synchronization: To strictly assess temporal alignment, we report AV-A (Audio-Video Alignment) distance computed via Synchformer [62]. For lip synchronization specifically, we report the SyncNet Confidence (SNC) score [63]. Additionally, global cross-modal alignment is measured via ImageBind (IB).
  • Music Evaluation: Musical output is assessed using the SongEval[64], focusing on Coherence, Musicality, Memorability, Clarity, and Naturalness.
@@
MaskGCT[48] TTS 0.713 0.773 2.26 2.40
E2-TTS[49] 0.706 0.713 2.49 1.91
F5-TTS[18] 0.664 0.750 1.89 1.53
ZipVoice[50] 0.697 0.751 1.70 1.40
M3-TTS[22] 0.604 0.621 1.36 1.31
CosyVoice2[28] 0.652 0.748 2.57 1.45
@@ 6.4.1 Ablation Studies
Superiority of Natural Noise. Most notably, our experiments highlight the distinct advantage of using naturally collected white noise. As shown in Figure 3, the natural noise strategy achieves the lowest WER among all settings, significantly outperforming the zero-vector and Gaussian variants. Specifically, compared to the worst-performing setting (WER 3.17%), the natural noise setting achieves a WER of 2.06%, representing a substantial relative improvement of 1.11%. While this strategy results in a slight compromise in syncnet confidence dropping from the optimal 6.12 to 5.94 (a decrease of only 0.18), it maintains robust speaker timbre consistency. We attribute the superiority of natural noise to the manifold hypothesis: naturally collected noise lies closer to the true distribution of the background noise present in the training data compared to synthetic Gaussian distributions or null vectors. Consequently, the speaker encoder maps natural noise to a more effective region in the embedding space, providing a more semantically meaningful "negative" reference for the flow matching process. Based on these findings, we adopt naturally collected white noise as the default negative speaker embedding for inference.
Timbre Injection Impact Integrating the timbre injection mechanism introduces additional complexity to the conditioning space, which theoretically poses a risk of interfering with the model’s standard generation capabilities (T2VA/TI2VA). To quantify this trade-off, we conducted a side-by-side preference test comparing the full MM-Sonate model (supporting reference audio) against a variant trained without the timbre injection on standard non-cloning tasks. As shown in Figure 5(b), while there is a slight preference for the specialized text-only model in some cases (indicated by the "Worse" portion), the majority of samples are rated as "Same" or "Better" (GSB score: 0.89 for ZH, 0.83 for EN). We argue that this marginal degradation in standard tasks is an acceptable cost. Unlike the baseline, the full MM-Sonate unlocks entirely new generation paradigms—specifically TA2VA/TIA2VA, enabling zero-shot voice cloning capabilities that are otherwise impossible.
@@ 6.4.2 Evaluation of Audio-Video Generation
Figure 5: Subjective evaluation results. (a) Pass rates for speaker consistency, confirming robust identity preservation in both single-speaker and complex dialogue scenarios. (b) Ablation study comparing the full MM-Sonate against a baseline on standard generation tasks (T2VA/TI2VA). Although adding the timbre module causes a slight performance trade-off, the high "Same" and "Better" rates indicate that basic generation quality remains robust, while the model gains the unique capability for zero-shot voice cloning (TA2VA/TIA2VA).
@@ 6.4.4 Qualitative Analysis
To visually substantiate the quantitative metrics reported above, we present a frame-level analysis of generated samples across diverse task configurations in Figure 6. The figure illustrates MM-Sonate’s "omni-modal" capability, handling inputs ranging from text-only instructions (T2VA) to complex combinations involving reference images and audio (TIA2VA). Crucially, the time-aligned mel-spectrograms reveal strict synchronization between visual dynamics and acoustic events. In speech and voice cloning scenarios (Panels a-d), we observe precise alignment between lip closures and plosive sounds, as well as distinct mouth openings corresponding to vowel peaks. Remarkably, we observe that our timbre injection mechanism generalizes beyond human speech. As shown in Panel (b), when conditioned on animal vocalizations (e.g., a dog’s bark or a goat’s bleat), MM-Sonate performs cross-species acoustic cloning. It captures the specific textural identity of the reference animal sound and synthesizes a corresponding video where the animal’s jaw movements and breathing patterns physically match the acoustic impulses. This suggests that the model has learned a generalized representation of "sound source identity" rather than being limited to human phonetics. For musical generation (Panels e-f), the model accurately captures the rhythmic structure, synchronizing sustained vocal notes with prolonged visual articulations. Furthermore, in sound effect generation (Panels g-h), high-energy spectral bursts coincide perfectly with physical motion events. These qualitative results confirm that MM-Sonate maintains high-fidelity cross-modal coherence regardless of the conditioning modality or generative domain.

### v2

@@ 6.3 Evaluation Metrics
  • Video Quality: We report Motion Score (MS) based on RAFT Teed and Deng (2020) optical flow to quantify dynamic realism. Aesthetic Score (AS) is a composite metric derived from MANIQA Yang et al. (2022), aesthetic-predictor-v2-5 discus0434 (2024), and Musiq Ke et al. (2021) to assess visual fidelity. Identity preservation is measured by ID Consistency (ID), computed as the mean DINOV3 Siméoni et al. (2025) feature similarity between reference images and generated frames.
  • Audio Quality: We measure the distributional distance using Fréchet Distance (FD) and KL Divergence on mel-spectrogram features extracted via PANNs Kong et al. (2020). Semantic alignment between audio and text is evaluated using the CLAP score Wu et al. (2023).
  • Speech & Voice Cloning: Speech intelligibility is quantified by the Word Error Rate (WER), derived from Whisper-large-v3 Radford et al. (2023) transcriptions. To assess speaker identity preservation, we employ specific metrics depending on the evaluation context. For the main comparative benchmark against state-of-the-art TTS models (Table 2), we follow the SeedTTSAnastassiou et al. (2024) evaluation protocol and report SIM-o, calculated using the WavLM-large Chen et al. (2022) speaker verification model. In our ablation studies regarding negative conditioning strategies (Fig. 4 and 5), we utilize the Resemblyzer encoder***https://github.com/resemble-ai/Resemblyzer to measure Speaker Consistency.
  • Audio-Visual Synchronization: To strictly assess temporal alignment, we report AV-A (Audio-Video Alignment) distance computed via Synchformer Iashin et al. (2024). For lip synchronization specifically, we report the SyncNet Confidence (SNC) score Chung and Zisserman (2016). Additionally, global cross-modal alignment is measured via ImageBind (IB).
  • Music Evaluation: Musical output is assessed using the SongEvalYao et al. (2025), focusing on Coherence, Musicality, Memorability, Clarity, and Naturalness.
@@
MaskGCTWang et al. (2024b) TTS 0.713 0.773 2.26 2.40
E2-TTSEskimez et al. (2024) 0.706 0.713 2.49 1.91
F5-TTSChen et al. (2024) 0.664 0.750 1.89 1.53
ZipVoiceZhu et al. (2025) 0.697 0.751 1.70 1.40
M3-TTSWang et al. (2025e) 0.604 0.621 1.36 1.31
CosyVoice2Du et al. (2024b) 0.652 0.748 2.57 1.45
@@ 6.4.1 Ablation Studies
Superiority of Natural Noise. Most notably, our experiments highlight the distinct advantage of using naturally collected white noise. As shown in Figure 4, the natural noise strategy achieves the lowest WER among all settings, significantly outperforming the zero-vector and Gaussian variants. Specifically, compared to the worst-performing setting (WER 3.17%), the natural noise setting achieves a WER of 2.06%, representing a substantial relative improvement of 1.11%. While this strategy results in a slight compromise in syncnet confidence dropping from the optimal 6.12 to 5.94 (a decrease of only 0.18), it maintains robust speaker timbre consistency. We attribute the superiority of natural noise to the manifold hypothesis: naturally collected noise lies closer to the true distribution of the background noise present in the training data compared to synthetic Gaussian distributions or null vectors. Consequently, the speaker encoder maps natural noise to a more effective region in the embedding space, providing a more semantically meaningful ”negative” reference for the flow matching process. Based on these findings, we adopt naturally collected white noise as the default negative speaker embedding for inference.
Timbre Injection Impact Integrating the timbre injection mechanism introduces additional complexity to the conditioning space, which theoretically poses a risk of interfering with the model’s standard generation capabilities (T2VA/TI2VA). To quantify this trade-off, we conducted a side-by-side preference test comparing the full MM-Sonate model (supporting reference audio) against a variant trained without the timbre injection on standard non-cloning tasks. As shown in Figure 6(b), while there is a slight preference for the specialized text-only model in some cases (indicated by the ”Worse” portion), the majority of samples are rated as ”Same” or ”Better” (GSB score: 0.89 for ZH, 0.83 for EN). We argue that this marginal degradation in standard tasks is an acceptable cost. Unlike the baseline, the full MM-Sonate unlocks entirely new generation paradigms—specifically TA2VA/TIA2VA, enabling zero-shot voice cloning capabilities that are otherwise impossible.
@@ 6.4.2 Evaluation of Audio-Video Generation
Figure 6: Subjective evaluation results. (a) Pass rates for speaker consistency, confirming robust identity preservation in both single-speaker and complex dialogue scenarios. (b) Ablation study comparing the full MM-Sonate against a baseline on standard generation tasks (T2VA/TI2VA). Although adding the timbre module causes a slight performance trade-off, the high ”Same” and ”Better” rates indicate that basic generation quality remains robust, while the model gains the unique capability for zero-shot voice cloning (TA2VA/TIA2VA).
@@ 6.4.4 Qualitative Analysis
To visually substantiate the quantitative metrics reported above, we present a frame-level analysis of generated samples across diverse task configurations in Figure 7. The figure illustrates MM-Sonate’s ”omni-modal” capability, handling inputs ranging from text-only instructions (T2VA) to complex combinations involving reference images and audio (TIA2VA). Crucially, the time-aligned mel-spectrograms reveal strict synchronization between visual dynamics and acoustic events. In speech and voice cloning scenarios (Panels a-d), we observe precise alignment between lip closures and plosive sounds, as well as distinct mouth openings corresponding to vowel peaks. Remarkably, we observe that our timbre injection mechanism generalizes beyond human speech. As shown in Panel (b), when conditioned on animal vocalizations (e.g., a dog’s bark or a goat’s bleat), MM-Sonate performs cross-species acoustic cloning. It captures the specific textural identity of the reference animal sound and synthesizes a corresponding video where the animal’s jaw movements and breathing patterns physically match the acoustic impulses. This suggests that the model has learned a generalized representation of ”sound source identity” rather than being limited to human phonetics. For musical generation (Panels e-f), the model accurately captures the rhythmic structure, synchronizing sustained vocal notes with prolonged visual articulations. Furthermore, in sound effect generation (Panels g-h), high-energy spectral bursts coincide perfectly with physical motion events. These qualitative results confirm that MM-Sonate maintains high-fidelity cross-modal coherence regardless of the conditioning modality or generative domain.
@@ 8 Conclusion
Acknowledgement
We acknowledge the contributions from (sorted by first name): Boyuan Jiang, Chen Zhang, Feng Deng, Jiahao Wang, Jingbin He, Jingke Li, Jingru Zhao, Junjie Yan, Liang Hou, Lingyu Zou, Ming Wen, Nan Li, Peihan Li, Pengfei Wan, Teng Ma, Xiaoyu Shi, Xijuan Zeng, Xin Tao, Xu Li, Yan Zhou, Yiran Wang, Yu Zhao, Yuan Gao, Yun Li, Yushen Chen, Yuzhe Liang, Zewen Song, Zhongliang Liu, Zihan Li, Zihao Ji, Ziyang Yuan, and Ziyu Zhang.